# The Lunar Gravitational-wave Antenna: Mission Studies and Science Case

### v1

@@ 1 Introduction
  • Hubble constant measurement. Only LGWA can detect GW signals from DWDs outside our galaxy. Only DWDs produce strong enough signals and are so ubiquitous in the local Universe that a large enough number can be detected with identified host galaxies to measure the Hubble constant accurately with the help of GW observations.
@@
The purpose of this paper is to provide a description of the LGWA science case and an outline of the scientific studies needed for mission planning. Instead, we do not enter into the details of the payload design and technological requirements, and of the possible deployment scenarios, which will be provided with a dedicated document.
@@ 2.2.2 Localization of GW sources in space and time
The localization of GW sources is an important aspect of GW observations. Unless one considers the multiband scenario described in section 2.2.4, LGWA is likely going to be an isolated detector whose localization capabilities are limited compared to detector networks. However, there are important exceptions. Solar-mass compact binaries consisting of white dwarfs, neutron stars and lighter black holes will be observed by LGWA for several weeks to months. The modulation of the signal’s amplitude and phase due to the Moon’s rotation and orbital motion provide information about the source’s location in the sky. We also point out that contrary to laser-interferometric detectors, LGWA carries out independent surface displacement measurements along two orthogonal horizontal directions, which also helps with the sky localization. Vice versa, being able to measure sky location accurately improves estimates of other parameters like distance and polarization. Such a scenario is specifically attractive for BNS observations, where the detection and localization of the source would become an early warning of a BNS merger in the band of terrestrial detectors, and electromagnetic observatories can prepare for the event [57, 58]. Also the disruption or merger of white dwarfs observable in the decihertz band would be accompanied by EM counterparts (see sections 3.2.1 and 3.3.2), and sky localization might be crucial for the detection of an EM signal.
@@
Since for LGWA the motion of the Moon around the Earth and around the Sun needs to be simulated together with its rotation, the natural choice for LGWA seems to be to use a solar-centered system. However, this is just convenient from a simulation perspective, and does not solve the problem. For these reasons, we do not present a detailed assessment of sky-localization and merger-time errors here, and leave it instead to a future study.
Heuristically, one may select the reference frame in order to get as close as possible to the expected result. This was done to analyze the localization of GW170817 with LGWA using the Fisher-matrix approximation [14], and it was shown that few-arcmin2 sky-localization errors can be obtained. Similar sky-localization capabilities should be expected for any GW signal that has similar SNR and can be observed with LGWA over months to years like galactic DWDs and solar-mass BBHs. The order of magnitude of this result is not dependent on our (still uncertain) Fisher-matrix analyses alone, but can also be estimated analytically under simplified assumptions, e.g., neglecting changing antenna patterns and correlations between sky-location parameters and other waveform parameters [60, eq. (52)].
@@ 2.2.3 Estimates of spin and mass parameters
Estimation of mass and spin parameters can be important for various reasons. Component masses together with the distance of compact binaries are the fundamental quantity to start a population study. Furthermore, spins can help to distinguish formation channels of compact binaries. The LGWA science case on population studies and the role of the various parameters are described in section 3.2.2. In this section, we provide information about the capability of LGWA to measure these intrinsic parameters. The method is based on the Fisher-matrix approximation. The waveforms were simulated with the approximant IMRPhenomXPHM [61], which can simulate precession and higher-order modes.
@@ 2.2.4 Multiband GW observations
  • Stellar Origin Binary Black Holes. These sources are the primary target of the LIGO and Virgo detectors, with a total of $\sim 90$ sources already observed [2]. Their multi-band potential is actively studied in the context of LISA and ground-based interferometers [69, 77, 78, 79, 80, 81, 82, 83, 84, 85]. For example, a binary Black Hole like GW150914 on a circular orbit, with source-frame masses of $36$ and $29$ solar masses [1], was emitting at $\sim 0.016\,\rm Hz$ $5\,\rm yrs$ before merger, well inside the LISA band. The same system $1$ day before merger emits at $\sim 0.26\,\rm Hz$, which is around the peak sensitivity of LGWA. This shows that Stellar-origin BBHs are natural candidates for joint detections even in all the three bands accessible to LISA, LGWA, and ground-based detectors. A particularly interesting possibility would be the measurement in LGWA of binary parameters hardly accessible to ground-based detectors, such as eccentricity.
@@
  • Binary Neutron Stars. A multiband detection of these systems would be a unique possibility opened by LGWA, since those are not accessible to LISA. For example, a Binary Neutron Star system with source-frame chirp mass $\mathcal{M}_{c}=1.188\,M_{\odot}$, corresponding to the value measured for GW170817 [91] was emitting at $\sim 0.05\rm Hz$, at the edge of the LISA band, more than 45 yrs before the merger. This timescale is clearly too long for the possibility of a joint detection. On the contrary, such a source emits at $\sim 0.27\,\rm Hz$, around the peak sensitivity of LGWA, 6 months before merger, being a perfect candidate for joint observations of LGWA and ground-based detectors, which can observe BNS systems at high rates [89]. Considering 10 years as a limit timescale for the possibility of a joint detection, a GW170817-like source could be observed from a starting frequency of $\sim 0.09\,\rm Hz$. We estimate the joint detection rate in LGWA and ET using the same procedure as for stellar-origin BBHs. For BNS systems, the uncertainty on the population model is much larger; in particular, the value of the local merger rate inferred from the GWTC–3 catalog, assuming a flat mass distribution, is $\mathcal{R}_{0}=105^{+190}_{-83}\,\rm Gpc^{-3}\,yr^{-1}$. Using the central value of $105\,\rm Gpc^{-3}\,yr^{-1}$, a flat mass distribution between $[1.1,2.5]\,\rm M_{\odot}$, and values for the Madau-Dickinson profile obtained as described in Appendix A of [89], we obtain $\sim 36$ sources per year within redshift $z\sim 0.1$, which corresponds to the LGWA horizon at $\sim\rm few\,M_{\odot}$. Over a period of $10\,\rm yrs$, we obtain 5 joint detections between LGWA and ET with $\rm SNR>8$. Given the large uncertainty on the local merger rate, we also compare to the results obtained using the upper and lower values of $295\,\rm Gpc^{-3}\,yr^{-1}$ and $22\,\rm Gpc^{-3}\,yr^{-1}$, that give 24 and 1 joint LGWA-ET detections respectively. We conclude that a joint detection is a possibility that should be expected in LGWA. The particularly long time spent by BNS systems in the LGWA band would allow in particular their precise localization prior to merger, which is a key need for potential multimessenger observations. Also, long-duration multi-band studies using LGWA and ground-based GW detectors [71, 92] can improve the estimation of parameters, in particular the tidal deformability [70]. This will lead to better constraints, ruling out many more theoretical EOS models [92].
@@
  • Supermassive Black Holes. These systems represent a target for joint LISA-LGWA detections. In particular, there is ample overlap in the mass and redshift range of the two observatories between $10^{3}-10^{6}\,M_{\odot}$ up to redshift $\sim 10$. For example, Ref. [93] finds between $\approx 25$ and $\approx 75$ events per year depending on the seed model. Most or all of them will be detected also by LGWA, which can in particular accumulate enough late inspiral-merger-ringdown SNR allowing a better localization of the sources, especially at high redshift. This is of paramount importance since a combined LISA-LGWA analysis might help pinpointing the source parameters with higher accuracy, in particular the source distance. In fact, although LISA can see MBHBs at $z>20$, for faint sources the distance estimate is rather poor and determining the actual high redshift nature of the source might be challenging [94, 95]. Moreover, LGWA will be sensitive to the merger and ringdown part of the signal, which for systems of $M<10^{4}$ at $z>5$ is not accessible to LISA.
@@ 2.3.4 Ground tilt response
Viscoelastic relaxation processes in the Moon interior affect the tidal response by introducing a frequency dependence of the tidal processes and a lag between each tidal harmonic and the corresponding response, which is directly related to dissipation of tidal energy into the Moon mantle [see, e.g., 127]. While the frequency dependent effects are generally small compared to the elastic response, precise observation of the amplitudes and phase lags at different tidal frequencies would provide an indirect probe into the rheology and the structure of the Moon interior [128, 129].
@@ 3.1.1 Seismic sources
Apollo seismic observation revealed that the Moon is still seismically active today. After about seven years of seismic observation on the Moon, more than 13000 lunar seismic events (known as moonquakes) were identified and cataloged [178]. The catalog has been constantly updated with state-of-the-art methods applied to the data [179, 180], and even today, new events are discovered from the data [e.g 181]. The Moon has different types of seismic sources, including the endogenic ones. This opened a new question of where and how these seismic events are excited. This is one of the key open questions in lunar seismology and was listed as one of the science goals by Artemis Science Definition Team [182].
When mapping the global scale events distribution for the Moon, the seismicity on the lunar farside is less constrained. Both, the Apollo seismic network and most of the detected moonquake sources, were located on the lunar nearside [132, 134]. A minimal number of earthquakes on the lunar farside were identified, and it remains uncertain whether this is a result of the Apollo seismic network’s limited coverage or if the farside is genuinely aseismic. Given the limited sensitivity of the Apollo seismometers, it is very likely that the network did not have, for the farside, the same detection capability as for the nearside, especially for small magnitude events that dominate the catalogue [e.g. 183]. On the other hand, global observation carried out after Apollo revealed that there is a clear dichotomy on the Moon, and there are significant differences in terms of internal structure between the lunar nearside and the farside [e.g. 184, 185]. These may result in variations of the seismicity rate and spatial distribution, and to elucidate this open question, new observation at different location from Apollo will be essential. The dichotomy on the Moon is one of the key question to understand the evolution of the Moon. Seismicity will inform us about the depth of the dichotomy, which serves as an important constraint on the crust/mantle formation.
@@ 3.2.1 Astrophysical explosions and matter effects on waveforms
Main contributors: Alexey Bobrick, Martina Toscani, David Vartanyan, Tsvi Piran, Jan Harms, Suvodip Mukherjee, Sourav Roy Chowdhury, Aayushi Doshi, Suraj N K, Trisha V
@@
Stellar mergers are a prominent example where GWs may inform us about physical processes. The LGWA has the unique capability to observe WD/WD mergers. Specifying the properties of two white dwarfs, it is currently not yet established with certainty what specific outcome their merger may produce [251, 252, 253]. The complexity arises because the nuclear detonation or deflagration conditions are extremely hard to resolve in simulations [254, 255, 256]. They also have a strong sensitivity to the geometry of the merger and the composition profiles, and yet, they play a decisive role in the outcome of the merger [257, 258]. In addition, these processes happen during the fairly complex gravitational interactions of the two stars [251, 259]. Moreover, following the onset of deflagration or detonation, the exact nucleosynthesis in the white dwarfs is also challenging to model [260, 261].
@@
There are multiple classes of astrophysical transients and outcomes known to be produced by double white dwarf mergers [262]. As an example, it is known that some double white dwarf binaries may detonate fully and lead to bright supernovae such as normal SNe Type Ia, 91bg-like SNe, SNe Type 91T, or Super-Chandrasekhar SNe ([263], see also Sec. 3.3.2). Similarly, some white dwarf binaries may experience a weak detonation and produce Ca-strong SNe, SNe Type Iax, 08ha-like SNe [263]. In some other binaries, the ignition may be even more gentle, leaving behind remnants such as another, more massive, white dwarf, an sdB star or an RCorBor star [251]. LGWA observations may be transformative for our understanding of such events. Identifying the connection between the progenitor white dwarf masses and the transient event they produce would dramatically improve our understanding and modelling efforts for these systems. Moreover, deci-Hertz GW observations may be the only way of probing what happened to the binary material in the first seconds following the detonation, in particular shedding light on the fate of the surviving object and the physics of explosion. The main caveat in this regard is the relatively low rate of double white dwarf mergers in the Galaxy and the local universe.
@@
Less explored avenues involve the physics of mergers of white dwarfs with neutron stars, stellar black holes [264, 265] and intermediate-mass black holes [266]. In all these cases, such mergers are expected to produce an optical transient event and are governed by a complex interplay of nuclear evolution, magnetohydrodynamics and radiative transfer, especially in the vicinity of the compact object, which is the dominant feedback source in such events [267, 268, 269].
Figure 26: Gravitational wave signal (solid blue line) produced as part of the 3-dimensional magnetohydrodynamic simulation of the merger of a 1.4 M NS with a 1 M carbon-oxygen WD companion at arbitrary distance of 50 kpc [270].
@@
(see [278] for relativistic corrections to this formula) and $\beta=R_{\rm t}/R_{\rm p}$ is the penetration factor, that is the ratio between the tidal radius and the stellar pericenter. In reality, the exact extent of the disruption depends on the specific characteristics and internal structure of the star itself [see, e.g., 279]. In general, we can categorize three distinct cases: when $\beta\gtrsim 1$ the star experiences full disruption; when $\beta\approx 1$ the star undergoes partial disruption, possibly leading to quasi-periodic eruptions (e.g., [280, 281]; see Sec.3.3.5, for details); and when $\beta\lesssim 1$ no disruption of the star occurs.
@@
Interactions in stably transferring binaries are also physically rich, and involve accretion disc physics, hydrodynamics, possible jet and outflow formation, and still poorly understood mechanisms of mass loss [288, 289, 290, 291]. With GW observations, it may be possible to constrain the mass transfer rates in these systems, this way at least indirectly constraining the nature of interactions [292]. Coupling such observations with optical, X-ray and radio data may be particularly constraining for understanding the interaction physics in such binaries [293, 294]. For the Galactic sources, the timescales in the inner accretion disc match the LGWA band and, therefore, for sufficiently massive accretion discs, such observations may potentially be important [295, 296].
Mergers involving bigger objects may appear too slow to produce a measurable signal in the LGWA band. However, the compact central cores of geometrically large objects may have much shorter and much more relevant timescales. For example, common envelope merger events of giants and compact stars that initially happen on year-long timescales have been recently shown to produce a measurable GW signal for LISA [297, 298, 269]. Similarly, in mergers of massive main sequence stars, which take place on tens of minutes timescales, the geometrically smaller cores may merge on the factor of tens timescales shorter [299]. Such mergers are common in the Galaxy. For similar reasons, mergers of white dwarfs with main sequence stars or white dwarfs with the cores of red giants or stars of the asymptotic giant branch may be visible.
Tidal interactions in highly eccentric systems may also have their unique GW signature. Such binaries form in hierarchical triple stellar systems [300, 301]. While the effective timescale of the periaston passage may be in tens of minutes timescale, the tidal oscillations induced in the tidally deformed stars may have a factor of several higher frequency harmonics. Potentially detecting such signals will be informative both of the tidal interactions and the triple stellar dynamics for systems containing tidally interacting stars.
@@
Another unique avenue is the detection of GW signature from the helium flash at the tip of the red-giant branch. Ignition of helium is expected to be off-center, have a spherically-asymmetric formation and happen on timescales of tens to hundreds of seconds [302, 303]. Due to the complexity of the onset of helium flash, little nucleo-hydrodynamic modelling has been done for it. Due to it happening in the very centres of giants, GWs, along with neutrinos, may be the only way of observationally constraining these events.

### v2

@@ 1 Introduction
  • Hubble constant measurement. Only LGWA can detect GW signals from DWDs beyond our galaxy and nearby dwarf galaxies. Only DWDs produce strong enough signals and are so ubiquitous in the local Universe that a large enough number can be detected with identified host galaxies to measure the Hubble constant accurately with the help of GW observations.
@@
The purpose of this paper is to provide a description of the LGWA science case and an outline of the scientific studies needed for mission planning. Instead, we do not enter into the details of the payload design and technological requirements, and of the possible deployment scenarios, which will be provided with a dedicated document. There are also important limitations to the analyses presented in this paper. First and foremost, except for selected cases, there are no accurate simulations available yet to estimate the sky-localization capabilities of LGWA alone or as part of a network with LISA-type detectors or terrestrial detectors. This problem has analytical as well as numerical-precision aspects that must be solved. Furthermore, we lack accurate waveform models of some signals where matter effects are important, and population models are generally less developed compared to most models used in science-case studies of terrestrial detectors and LISA.
@@ 2.2.2 Localization of GW sources in space and time
The localization of GW sources is an important aspect of GW observations. Unless one considers the multiband scenario described in section 2.2.4, LGWA is likely going to be an isolated detector whose localization capabilities are limited compared to detector networks. However, there are important exceptions. Solar-mass compact binaries consisting of white dwarfs, neutron stars and lighter black holes will be observed by LGWA for several weeks to years. The modulation of the signal’s amplitude and phase due to the Moon’s rotation and orbital motion provide information about the source’s location in the sky. We also point out that contrary to laser-interferometric detectors, LGWA carries out independent surface displacement measurements along two orthogonal horizontal directions, which also helps with the sky localization. Vice versa, being able to measure sky location accurately improves estimates of other parameters like distance and polarization. Such a scenario is specifically attractive for BNS observations, where the detection and localization of the source would become an early warning of a BNS merger in the band of terrestrial detectors, and electromagnetic observatories can prepare for the event [56, 57]. Also the disruption or merger of white dwarfs observable in the decihertz band would be accompanied by EM counterparts (see sections 3.2.1 and 3.3.2), and sky localization might be crucial for the detection of an EM signal.
@@
Figure 9: Sky-localization estimates based on [59, eqs. (51),(52)] coming from the orbital motion of the Moon (around the Earth) and the Earth (around the Sun). The bands represent a range of frequencies from $f=0.02$ Hz (worse localization) to 0.5 Hz (better localization) with the solid lines representing 0.1 Hz. The signal SNR increases with the square-root of the observation time reaching a value of 30 after one year.
A correction can however be applied for selected signals by expressing the detector positions with respect to a reference point that minimizes coordinate artifacts. This was done to analyze the localization of GW170817 with LGWA using the Fisher-matrix approximation [14], and it was shown that few-arcmin2 sky-localization errors can be obtained. The order of magnitude of this result can also be estimated analytically under simplified assumptions, e.g., neglecting changing antenna patterns and assuming a fixed signal frequency [59, eqs. (51),(52)]. Results from the analytical expressions are shown in figure 9.
Accordingly, the signal modulation caused by the orbital motion around the Sun leads to the dominant sky-localization capabilities of LGWA compared to the contribution of the Moon’s orbital motion. Sky localization improves quadratically with the SNR and signal frequency.
@@ 2.2.3 Estimates of spin and mass parameters
Estimation of mass and spin parameters can be important for various reasons. Component masses together with the distance of compact binaries are the fundamental quantity to start a population study. Furthermore, spins can help to distinguish formation channels of compact binaries. The LGWA science case on population studies and the role of the various parameters are described in section 3.2.2. In this section, we provide information about the capability of LGWA to measure these intrinsic parameters. The method is based on the Fisher-matrix approximation. The waveforms were simulated with the approximant IMRPhenomXPHM [60], which can simulate precession and higher-order modes. The Fisher-matrix method leads to accurate error estimates whenever the SNR is high enough and there is no degeneracy of the waveform model involving the analyzed parameters. These conditions can be assessed post factum when finding that the parameter-estimation errors are small. Large parameter errors point to potential issues typically with model degeneracies, which can be cured by using prior distributions as well [61], or by switching to a full Bayesian analysis tool without assuming a Gaussian likelihood [62]. This means that the errors presented in this section for mass parameters are a good indication of what a full Bayesian analysis would obtain, while errors of the spin parameters might deviate significantly from full-Bayesian calculations when the error approaches unity. Error estimates from a full Bayesian analysis are typically significantly smaller in this case due to the help of (even uniform) prior distributions [61].
@@ 2.2.4 Multiband GW observations
  • Stellar Origin Binary Black Holes. These sources are the primary target of the LIGO and Virgo detectors, with a total of $\sim 90$ sources already observed [2]. Their multi-band potential is actively studied in the context of LISA and ground-based interferometers [70, 78, 79, 80, 81, 82, 83, 84, 85, 86]. For example, a binary Black Hole like GW150914 on a circular orbit, with source-frame masses of $36$ and $29$ solar masses [1], was emitting at $\sim 0.016\,\rm Hz$ $5\,\rm yrs$ before merger, well inside the LISA band. The same system $1$ day before merger emits at $\sim 0.26\,\rm Hz$, which is around the peak sensitivity of LGWA. This shows that Stellar-origin BBHs are natural candidates for joint detections even in all the three bands accessible to LISA, LGWA, and ground-based detectors. A particularly interesting possibility would be the measurement in LGWA of binary parameters hardly accessible to ground-based detectors, such as eccentricity. Detecting signals that cross a wide band from low to high frequencies would furthermore strengthen the constraints on GW speed propagation [87], tests of GR [88, 89, 90, 91] and the whole set of source parameters [70]. In particular, the accurate sky localization can lead to the unique possibility of a pre–alert of telescopes for the detection of potential counterparts, which are possible in some astrophysical environments; knowledge of the eccentricity can be combined with spin measurements by ground–based detectors to constrain BBH formation channels; precise knowledge of the inspiral part of the signal will enable unprecedented accuracy in constraining the merger physics.
@@
  • Binary Neutron Stars. A multiband detection of these systems would be a unique possibility opened by LGWA, since those are not accessible to LISA. For example, a Binary Neutron Star system with source-frame chirp mass $\mathcal{M}_{c}=1.188\,M_{\odot}$, corresponding to the value measured for GW170817 [97] was emitting at $\sim 0.05\rm Hz$, at the edge of the LISA band, more than 45 yrs before the merger. This timescale is clearly too long for the possibility of a joint detection. On the contrary, such a source emits at $\sim 0.27\,\rm Hz$, around the peak sensitivity of LGWA, 6 months before merger, being a perfect candidate for joint observations of LGWA and ground-based detectors, which can observe BNS systems at high rates [95]. Considering 10 years as a limit timescale for the possibility of a joint detection, a GW170817-like source could be observed from a starting frequency of $\sim 0.09\,\rm Hz$. We estimate the joint detection rate in LGWA and ET using the same procedure as for stellar-origin BBHs. For BNS systems, the uncertainty on the population model is much larger; in particular, the value of the local merger rate inferred from the GWTC–3 catalog, assuming a flat mass distribution, is $\mathcal{R}_{0}=105^{+190}_{-83}\,\rm Gpc^{-3}\,yr^{-1}$. Using the central value of $105\,\rm Gpc^{-3}\,yr^{-1}$, a flat mass distribution between $[1.1,2.5]\,\rm M_{\odot}$, and values for the Madau-Dickinson profile obtained as described in Appendix A of [95], we obtain $\sim 36$ sources per year within redshift $z\sim 0.1$, which corresponds to the LGWA horizon at $\sim\rm few\,M_{\odot}$. Over a period of $10\,\rm yrs$, we obtain 5 joint detections between LGWA and ET with $\rm SNR>8$. Given the large uncertainty on the local merger rate, we also compare with the results obtained using the upper and lower rates values of $295\,\rm Gpc^{-3}\,yr^{-1}$ and $22\,\rm Gpc^{-3}\,yr^{-1}$, which give 24 and 1 joint LGWA-ET detections respectively. We conclude that a joint detection is a possibility that should be expected in LGWA. The particularly long time spent by BNS systems in the LGWA band would enable their precise localization prior to merger. Combined with the early detection by ground-based 3G detectors, the sky patch to scan for EM signals would further shrink making use of the long baseline between Earth and Moon. There are prediction of radio and X-ray counterparts produced well before the merger of two neutron stars [98]. Early warnings from LGWA would make it possible to prepare for the observation of these faints signals. Also, long-duration multi-band studies using LGWA and ground-based GW detectors [72, 99] can improve the estimation of parameters, in particular the tidal deformability [71]. This will lead to better constraints, ruling out many more theoretical EOS models [99].
@@
  • Supermassive Black Holes. These systems represent a target for joint LISA-LGWA detections. In particular, there is ample overlap in the mass and redshift range of the two observatories between $10^{3}-10^{6}\,M_{\odot}$ up to redshift $\sim 10$. For example, Ref. [Bonetti:2018tpf] finds between $\approx 25$ and $\approx 75$ events per year depending on the seed model. Most or all of them will be detected also by LGWA, which can in particular accumulate enough late inspiral-merger-ringdown SNR allowing a better localization of the sources, especially at high redshift. This is of paramount importance since a combined LISA-LGWA analysis might help pinpointing the source parameters with higher accuracy, in particular the source distance. In fact, although LISA can see MBHBs at $z>20$, for faint sources the distance estimate is rather poor and determining the actual high redshift nature of the source might be challenging [100, 101]. Moreover, LGWA will be sensitive to the merger and ringdown part of the signal, which for systems of $M<10^{4}$ at $z>5$ is not accessible to LISA.
@@ 2.3.4 Ground tilt response
Viscoelastic relaxation processes in the Moon interior affect the tidal response by introducing a frequency dependence of the tidal processes and a lag between each tidal harmonic and the corresponding response, which is directly related to dissipation of tidal energy into the Moon mantle [see, e.g., 133]. While the frequency dependent effects are generally small compared to the elastic response, precise observation of the amplitudes and phase lags at different tidal frequencies would provide an indirect probe into the rheology and the structure of the Moon interior [134, 135].
@@ 3.1.1 Seismic sources
Apollo seismic observation revealed that the Moon is still seismically active today. After about seven years of seismic observation on the Moon, more than 13000 lunar seismic events (known as moonquakes) were identified and cataloged [183]. The catalog has been constantly updated with state-of-the-art methods applied to the data [184, 142], and even today, new events are discovered from the data [e.g 185]. The Moon has different types of seismic sources, including the endogenic ones. This opened a new question of where and how these seismic events are excited. This is one of the key open questions in lunar seismology and was listed as one of the science goals by Artemis Science Definition Team [186].
When mapping the global scale events distribution for the Moon, the seismicity on the lunar farside is less constrained. Both, the Apollo seismic network and most of the detected moonquake sources, were located on the lunar nearside [138, 140]. A minimal number of earthquakes on the lunar farside were identified, and it remains uncertain whether this is a result of the Apollo seismic network’s limited coverage or if the farside is genuinely aseismic. Given the limited sensitivity of the Apollo seismometers, it is very likely that the network did not have, for the farside, the same detection capability as for the nearside, especially for small magnitude events that dominate the catalogue [e.g. 187]. On the other hand, global observation carried out after Apollo revealed that there is a clear dichotomy on the Moon, and there are significant differences in terms of internal structure between the lunar nearside and the farside [e.g. 188, 189]. These may result in variations of the seismicity rate and spatial distribution, and to elucidate this open question, new observation at different location from Apollo will be essential. The dichotomy on the Moon is one of the key question to understand the evolution of the Moon. Seismicity will inform us about the depth of the dichotomy, which serves as an important constraint on the crust/mantle formation.
@@ 3.2.1 Astrophysical explosions and matter effects on waveforms
Main contributors: Alexey Bobrick, Martina Toscani, David Vartanyan, Tsvi Piran, Jan Harms, Suvodip Mukherjee, Sourav Roy Chowdhury, Aayushi Doshi, Suraj N K, Trisha V, Kiranjyot Gill
The decihertz band harbors signals from a multitude of astrophysical processes not accessible to the currently operating GW observatories. Their detection would transform the study of these objects. However, the GW signatures of many astrophysical objects are either uncertain or unknown, making it harder to estimate their expected rates and the information one might extract from the signals. If no significant improvement of the models is made, it would be harder to detect and analyze the signals. With this caveat in mind, however, thanks to the diversity of the potentially observable sources, some will likely be observed during the mission lifetime, possibly in conjunction with EM observations, this way significantly impacting the respective fields. We also note that the rates of individual detectable astrophysical events are sufficiently low so that no complication from overlapping waveforms is expected. Below, we review the most likely such sources.
@@
Stellar mergers are a prominent example where GWs may inform us about physical processes. The LGWA has the unique capability to observe WD/WD mergers. In the Galaxy, these mergers happen at a rate of about $3$ times per century [255]. However, specifying the properties of two white dwarfs, it is currently not yet established with certainty what specific outcome their merger may produce [256, 257, 258]. The complexity arises because the nuclear detonation or deflagration conditions are extremely hard to resolve in simulations [259, 260, 261]. They also have a strong sensitivity to the geometry of the merger and the composition profiles, and yet, they play a decisive role in the outcome of the merger [262, 263]. In addition, these processes happen during the fairly complex gravitational interactions of the two stars [256, 264]. Moreover, following the onset of deflagration or detonation, the exact nucleosynthesis in the white dwarfs is also challenging to model [265, 266].
@@
There are multiple classes of astrophysical transients and outcomes known to be produced by double white dwarf mergers [267]. As an example, it is known that some double white dwarf binaries may detonate fully and lead to bright supernovae such as normal SNe Type Ia, 91bg-like SNe, SNe Type 91T, or Super-Chandrasekhar SNe ([268], see also Sec. 3.3.2). They happen at a Galactic rate of approximately $6$ times less frequent than all DWD mergers, i.e., once per about two centuries [255]. However, due to their energetic nature, their GW signal is likely to be brighter [269]. Therefore, while it still remains to be studied, LGWA may potentially see such events extragalactically and hence more frequently. Similarly, some white dwarf binaries may experience a weak detonation and produce Ca-strong SNe, SNe Type Iax, 08ha-like SNe [268], with their deci-Hertz band GW signal being uncertain. In some other binaries, the ignition may be even more gentle, leaving behind remnants such as another, more massive, white dwarf, an sdB star or an RCorBor star [256]. The latter has been recently constrained to happen between once per 200 – 1000 years in the Galaxy[270]. LGWA observations may be transformative for our understanding of such events. Identifying the connection between the progenitor white dwarf masses and the transient event they produce would dramatically improve our understanding and modelling efforts for these systems. Moreover, deci-Hertz GW observations may be the only way of probing what happened to the binary material in the first seconds following the detonation, in particular shedding light on the fate of the surviving object and the physics of explosion. The main caveat in this regard is the relatively low rate of double white dwarf mergers in the Galaxy and the local universe, as well as the uncertainties related to their GW signal.
@@
Less explored avenues involve the physics of mergers of white dwarfs with neutron stars, stellar black holes [271, 272] and intermediate-mass black holes [273]. WD-NS mergers, in particular, happen at Galactic rates of about once per few thousand years [274, 275], but would be observed more frequently if extragalactic events were detectable. Specifically for extragalactic events, the EM signal for the event may allow one to search for the GW counterpart using waveform templates, potentially enhancing GW detectability. In all these cases, such mergers are expected to produce an optical transient event and are governed by a complex interplay of nuclear evolution, magnetohydrodynamics and radiative transfer, especially in the vicinity of the compact object, which is the dominant feedback source in such events [276, 277, 278].
Figure 27: Top panel: 3-dimensional magnetohydrodynamic simulation of the merger between a 1.4 Mdirect-$\odot$ neutron star and a 1 Mdirect-$\odot$ carbon-oxygen white dwarf companion, as presented in [279]. Panels (a)–(j) depict density slices in the orbital plane, starting from the beginning of the simulation (panel a) and continuing until t=242 s (panel j). Bottom panel: Corresponding gravitational wave signal assuming a distance of 50 kpc. The left plot shows the strain versus time, while the right plot presents the characteristic strain versus gravitational wave frequency.
@@
(see [287] for relativistic corrections to this formula) and $\beta=R_{\rm t}/R_{\rm p}$ is the penetration factor, that is the ratio between the tidal radius and the stellar pericenter. In reality, the exact extent of the disruption depends on the specific characteristics and internal structure of the star itself [see, e.g., 288]. In general, we can categorize three distinct cases: when $\beta\gtrsim 1$ the star experiences full disruption; when $\beta\approx 1$ the star undergoes partial disruption, possibly leading to quasi-periodic eruptions (e.g., [289, 290]; see Sec.3.3.5, for details); and when $\beta\lesssim 1$ no disruption of the star occurs.
@@
Interactions in stably transferring binaries are also physically rich, and involve accretion disc physics, hydrodynamics, possible jet and outflow formation, and still poorly understood mechanisms of mass loss [297, 298, 299, 300]. With GW observations, it may be possible to constrain the mass transfer rates in these systems, this way at least indirectly constraining the nature of interactions [301]. Coupling such observations with optical, X-ray and radio data may be particularly constraining for understanding the interaction physics in such binaries [302, 303]. For the Galactic sources, the timescales in the inner accretion disc match the LGWA band and, therefore, for sufficiently massive accretion discs, such observations may potentially be important [304, 305]. However, both the GW signal and the rates of detectable mergers need to be studied in detail.
Mergers involving bigger objects may appear too slow to produce a measurable signal in the LGWA band. However, the compact central cores of geometrically large objects may have much shorter and much more relevant timescales. For example, common envelope merger events of giants and compact stars that initially happen on year-long timescales have been recently shown to produce a measurable GW signal for LISA [306, 307, 278]. Similarly, in mergers of massive main sequence stars, which take place on tens of minutes timescales, the geometrically smaller cores may merge on the factor of tens timescales shorter [308]. Such mergers are common in the Galaxy and, for example, Renzo et al. [307] estimate that between $0.1$ and $100$ sources are detectable by LISA during its mission lifetime, implying comparable rates for LGWA. For similar reasons, mergers of white dwarfs with main sequence stars or white dwarfs with the cores of red giants or stars of the asymptotic giant branch may be visible, although modelling of their GW signatures and LGWA detection rates is still necessary. Generally, we envision that GW signatures of all types of stellar mergers mentioned here need to be studied in detail.
Tidal interactions in highly eccentric systems may also have their unique GW signature. Such binaries form in hierarchical triple stellar systems [309, 310]. While the effective timescale of the periaston passage may be in tens of minutes timescale, the tidal oscillations induced in the tidally deformed stars may have a factor of several higher frequency harmonics. Potentially detecting such signals will be informative both of the tidal interactions and the triple stellar dynamics for systems containing tidally interacting stars. However, the GW signatures of these systems also remain unknown so far.
@@
Another unique avenue is the detection of GW signature from the helium flash at the tip of the red-giant branch. The ignition of helium is expected to be off-center, exhibit non-spherical formation, and occur over timescales of tens to hundreds of seconds [311, 312]. Due to the complexity of the helium flash onset, limited attempts have been made thus far at nucleo-hydrodynamic modeling [313, 314]. Given these flashes occur deep within the center, GWs and neutrino observations may be the only way to constrain these events. While the Galactic rates of these events can be easily estimated as the formation rate of the red giant stars in the Galaxy [315], thus being approximately once per decade, their GW strain amplitudes are far from being understood, necessitating further research.